# Benchmarking Spiking Neural Network Learning Methods with Varying Locality

### v1

@@ III-D Centered Kernel Alignment
where, I is the identity matrix and O is a matrix consisting of all 1’s. Centering matrix C can be computed by . HSIC score of 0 means independence between two variables. CKA normalizes this into a similarity index from 0 to 1, where higher values imply greater similarity. We adhere to the use of an unbiased estimator for calculating HSIC across mini-batches [30, 43, 44]. CKA metric will also be used in later sections to empirically explain the performance differences of various SNN learning methods in terms of accuracy and robustness.
@@ IV-A Experimental Setup
Experiments conducted in this study includes two datasets for investigation of the performance of various training methods. The Neuromorphic-MNIST (N-MNIST) dataset is a neuromorphic adaptation of the handwritten digits classification task [45]. In contrast to the MNIST dataset [46], which consists of static images, N-MNIST encapsulates changes in pixel values as events unfold over time. N-MNIST is specifically designed for the evaluation of neuromorphic algorithms and models [47, 19]. Another dataset used in this study is the TIMIT dataset, which is a speech and audio processing sequence-to-sequence task [48, 13, 15]. Although both N-MNIST and TIMIT datasets involve the temporal dimension, data in TIMIT exhibits a higher degree of dependency between time steps.
Performance and robustness measurements are conducted for two types of SNN architectural designs. One is a feed-forward SNN (FF) and another is a recurrent SNN (REC) architecture, which introduces additional weights to the linear layers in FF. To classify N-MNIST dataset, both FF and REC SNNs consisting of two hidden layers of 120 and 84 neurons are used. The TIMIT classification task utilizes FF and REC SNNs with one hidden layer of 400 neurons. Notably, due to the configuration of e-prop, the output layer does not contain an explicit recurrent weight connection.
@@ IV-B Performance Evaluation
Training | N-MNIST | TIMIT | ||
Methods | REC | FF | REC | FF |
BPTT | ||||
e-prop | ||||
DECOLLE | ||||
This architecture trained on SMNIST dataset, uses three hidden layers with 64, 256, and 256 neurons [14]. |
This architecture uses three hidden layers with 512 neurons [31]. |
@@
This architecture uses one hidden layer with 800 neurons in an LSNN architecture [13], which is an LSTM-equivalent architecture in the spiking domain. In general, LSNN has better performance compared to Recurrent SNN. The performance is obtained from [15]. |
@@
This architecture uses one hidden layer with 800 or 400 neurons using an LSNN architecture [15]. |
@@
With 1 hidden layer of 1200 neurons, a feed-forward neural network trained with e-prop achieves ; With 3 hidden layers, each of which consists of 900 neurons, a feed-forward neural network trained with e-prop achieves accuracy [15]. |
@@
This architecture consists of 3 convolutional layers with 64, 128, and 128 channels respectively [19]. |
@@
Table I presents the outcomes of previous studies that have benchmarked performance of SNNs across two distinct types of classification tasks. These results show promising performance of SNNs in several classification tasks in comparison to state-of-the-art ANNs. However, due to the lack of consistency in neural network architecture designs and variations in training sets, direct comparison between different training methods is infeasible. To address this problem, we have conducted a series of benchmark tests with consistent neural network architectures and training datasets. This work aims to offer a more comprehensive and insightful assessment of the performance of local learning methods and the impact of explicit recurrence on SNNs.
Training | N-MNIST | TIMIT | ||
Methods | FF | REC | FF | REC |
BPTT | 97.12% | 97.77% | 52.31% | 56.93% |
e-prop | 96.80% | 96.44% | 51.12% | 54.99% |
DECOLLE | 93.61% | 90.52% | 42.30% | 45.70% |
@@
In our investigations, we have observed that the introduction of recurrent weights in the REC architecture yields significant performance improvements when classifying the TIMIT dataset, in contrast to the FF architecture. Table II shows SNN performance across various training methods for TIMIT and N-MNIST classification. When trained with BPTT, the REC architecture exhibits a notable enhancement in accuracy when compared to the FF architecture. Similar accuracy improvements are also observed in the case of e-prop and DECOLLE. Benefits of recurrent weights are not evident in the N-MNIST classification task. Here, the performance of SNNs trained with BPTT and e-prop displays only slight differences between the REC module and the FF module. Interestingly, the explicit inclusion of recurrent weights leads to a performance degradation in SNNs trained with DECOLLE for the N-MNIST dataset. Further investigation is required to understand the underlying reason for this degradation.
@@
Additionally, a more local learning method slightly degrades the performance of SNNs, which indicates a trade-off between biological plausibility and performance. In this study, the locality of learning method is considered to increase from BPTT (global learning method) to e-prop and DECOLLE successively. In N-MNIST classification, a slight accuracy difference is observed between BPTT and e-prop in both REC and FF modules. DECOLLE exhibits an accuracy reduction in both architectures compared with e-prop. A similar trend is observed for the TIMIT classification task.
@@ IV-C Representational Differences
CKA is a technique used to measure the similarity between two sets of data, and it has been increasingly applied to understand the internal representation in neural networks. Pioneering work by Cortes et al. introduced CKA as an alignment measure for kernels in the context of support vector machines [49]. Later, Kornblith et al. adapted CKA for deep learning, using it to compare layers across different neural network models, which provided insights into which layers of networks learn similar features [41]. Williams et al. expanded this analysis to study the similarity between representations learned by convolutional neural networks [50]. Li et al. investigated how SNNs differ from traditional ANNs in learning representations, using CKA to analyze both spatial and temporal dimensions [30]. In this study, CKA is employed as a metric to assess the similarity of representations across layers that are learned using various methods with differing degrees of local learning. Consistent with prior research, the CKA is averaged over 10 experiments with different initializations, each of which is conducted by passing 4096 instances from the testing set of N-MNIST dataset through the corresponding pre-trained model [30, 43]. We also compare the similarity of learnt representations of the various training methods with ANN-SNN conversion where the SNN is trained using conventional ANN training methods [51, 52].
@@ IV-D Robustness under Attack
Epsilon | 0 | 0.001 | 0.005 | 0.01 | 0.02 | 0.05 |
Training Methods | FF | |||||
BPTT | 97.12% | 83.82% | 16.78% | 4.34% | 1.22% | 0.73% |
e-prop | 96.80% | 89.90% | 41.27% | 12.33% | 3.27% | 2.87% |
DECOLLE | 93.61% | 83.48% | 45.11% | 28.57% | 21.29% | 17.79% |
Training Methods | REC | |||||
BPTT | 97.77% | 94.66% | 80.76% | 58.73% | 27.69% | 18.55% |
e-prop | 96.44% | 93.04% | 78.28% | 60.57% | 35.07% | 18.25% |
DECOLLE | 90.52% | 86.76% | 84.14% | 72.81% | 47.56% | 23.75% |
@@
DECOLLE | 90.21% | 0.86% | 87.41% | 27.75% | 81.55% | 88.49% | 81.29% | 88.96% | 80.81% | 91.02 % |
Specifically, the Fast Gradient Sign Method (FGSM) attack leverages the sign of gradients to create adversarial instances that maximize the loss of a neural network, misleading it to make incorrect classification [53]. However, due to the non-differentiable nature of SNNs, the direct calculation of gradients is not applicable. To solve this problem, previous work has proposed an effective framework which involves crafting an adversarial example for an SNN by creating its ANN counterpart and then calculating gradients for the corresponding ANN [54]. These gradients are subsequently utilized to construct an adversarial example for the SNN. We adopt this variation of FGSM attack to compare the robustness of various training methods and to investigate the impact of recurrent weights on robustness.
@@
Table III reports the performance of different training methods when subjected to FGSM attacks, where the perturbation level, denoted as “Epsilon”, varies across a range from 0.001 to 0.05. The results of our benchmarking analysis for backdoor attacks are illustrated in Table IV. We mainly focus on poison rates in realistic settings below 50% as shown in prior works [55, 56, 57, 35]. Both attacks are evaluated on the N-MNIST dataset.
Enhancement of the model’s robustness is observed with the inclusion of explicit recurrent weights. Under FGSM attacks, compared to FF models trained by the same methods and subjected to the same “Epsilon” values, the REC model retains more accuracy than the FF model. Besides, the robustness advantage of explicit recurrence is also observed under backdoor attack. Specifically, under more realistic settings with poison rates at 10% and 30%, we observe that the REC model consistently exhibits a lower ASR compared to its corresponding FF model, in all three training methods.
@@ IV-E CKA Measurements to Explain Robustness
It is observed that the addition of recurrent weights contributes to better preservation of the representations learned at each layer across these training methods. In these experiments, positive similarity differences are obtained at all perturbation levels, revealing that REC SNNs maintain larger representational similarity (implying more robust representations learned in the presence of attacks) compared to FF SNNs between clean and adversarial data. The results also reveal that global learning methods exhibit a greater enhancement in robustness with recurrent connections, as the similarity differences are larger in BPTT than in e-prop and DECOLLE. The trend matches the data provided in Table III. Through this analysis, we provide evidence that the inclusion of recurrent weights enhances the robustness of neuromorphic architectures.
@@ V Conclusions
In this work, learning methods with varying levels of locality are compared in terms of performance and robustness. To understand the performance disparity of these methods, representational differences are examined through the CKA metric. SNN learning methods with a higher degree of locality are shown to have performance degradation but they showed better robustness under FGSM adversarial attacks. However, the robustness advantages of local learning methods are not observed for non-gradient-based attacks like backdoor attacks. Additionally, we investigate the contribution of additional recurrent weights on the performance of SNNs. Measurements on the relative importance of recurrent weights are performed using Fisher Information. We illustrated that recurrent weights not only have positive influences on the performance of SNNs but also augment the robustness of SNNs. These observations are further substantiated by the fact that recurrent weights and linear weights exhibit comparable informative significance. In general, this work provides insights into the selection of learning methods for SNNs and offers a benchmarking analysis for future SNN designers. Future work can consider extension of such benchmarking analysis to more complex architectures including convolutional layers for image recognition tasks or attention layers for natural language processing tasks [16].
@@ Acknowledgments
This material is based upon work supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number #DE-SC0021562, the National Science Foundation grants CCF #1955815, CAREER # 2337646 and EFRI BRAID #2318101 and by Oracle Cloud credits and related resources provided by the Oracle for Research program.

### v2

@@ III-E Centered Kernel Alignment (CKA)
where, is the identity matrix and is a matrix consisting of all 1’s. HSIC score of 0 means independence between two variables. CKA normalizes this into a similarity index from 0 to 1, where higher values imply greater similarity. We adhere to the use of an unbiased estimator for calculating HSIC across mini-batches [29, 45, 46]. CKA metric will also be used in later sections to empirically explain the performance differences of various SNN learning methods in terms of accuracy and robustness.
@@ IV-A Experimental Setup
Experiments conducted in this study include Neuromorphic-MNIST (N-MNIST), DVS Gesture, and TIMIT for the investigation of the performance of various training methods.
N-MNIST: The N-MNIST dataset is a neuromorphic adaptation of the handwritten digits classification task [47]. In contrast to the MNIST dataset [48], which consists of static images, N-MNIST encapsulates changes in pixel values as events unfold over time. N-MNIST is specifically designed for the evaluation of neuromorphic algorithms and models [49, 19].
DVS Gesture: The DVS Gesture dataset [50] contains data from 29 subjects under three lighting conditions. A time step of 60 is used to evaluate performance [30].
TIMIT: The TIMIT dataset is a speech and audio processing sequence-to-sequence task [51, 13, 15]. Although N-MNIST, DVS Gesture, and TIMIT datasets all involve the temporal dimension, data in TIMIT exhibit a higher degree of dependency between time steps.
It is worth mentioning here that the e-prop method is currently limited to linear layers [15]. To ensure an equitable comparison between training methods without performance degradation, we therefore restrict our main experiments to fully connected architectures. To assess the generalizability of our conclusions, we further evaluate convolutional architectures trained with BPTT and DECOLLE in Appendix A. In this work, performance and robustness measurements are conducted for two types of SNN architectural designs. One is a feed-forward SNN (FF) and another is a recurrent SNN (REC) architecture, which introduces additional weights in the linear layers in FF. To classify the N-MNIST dataset, both FF and REC SNNs consisting of two hidden layers of 120 and 84 neurons are used. The DVS Gesture dataset is classified by FF and REC SNNs with one hidden layer of 512 neurons. The TIMIT classification task utilizes FF and REC SNNs with one hidden layer of 400 neurons. Notably, due to the configuration of e-prop, the output layer does not contain an explicit recurrent weight connection.
To ensure a fair comparison between the training methods, the hyperparameters of the SNN models trained with different paradigms were optimized for the best performance (Appendix B). The experiments were implemented using the snnTorch and PyTorch libraries in Python and conducted on an Nvidia RTX 2080 Ti GPU with 11GB of memory.
@@ IV-B Performance Evaluation
Training | N-MNIST | DVS Gesture | TIMIT | |||
Methods | FF | REC | FF | REC | FF | REC |
BPTT | ||||||
e-prop | ||||||
DECOLLE | ||||||
aThis architecture uses one hidden layer with 512 neurons[30]. | ||||||
@@
bThis architecture uses 3 hidden layers each with 900 neurons in an LSNN architecture [13], which is an LSTM-equivalent architecture in the spiking domain. In general, LSNN has better performance compared to recurrent SNN. The performance is obtained from [15]. | ||||||
@@
cThis architecture uses one hidden layer with 400 neurons using an LSNN architecture [15]. | ||||||
@@
dWith 1 hidden layer of 1200 neurons, a feed-forward LSNN trained with e-prop achieves accuracy; With 3 hidden layers, each of which consists of 900 neurons, a feed-forward LSNN trained with e-prop achieves accuracy [15]. | ||||||
eThis architecture uses one hidden layer with 400 neurons using an LSNN architecture [15]. | ||||||
@@
fThis architecture consists of 3 convolutional layers with 64, 128, and 128 channels respectively [19]. |
@@
Table II presents the outcomes of previous studies that have benchmarked the performance of SNNs. These results show promising performance of SNNs in several classification tasks in comparison to state-of-the-art ANNs. However, due to the lack of consistency in neural network architecture designs and variations in training sets, direct comparison between different training methods is infeasible. To address this problem, we have conducted a series of benchmark tests with consistent neural network architectures and training datasets. This work aims to offer a more comprehensive and insightful assessment of the performance of local learning methods and the impact of explicit recurrence on SNNs.
Training | N-MNIST | DVS Gesture | TIMIT | |||
Methods | FF | REC | FF | REC | FF | REC |
BPTT | 97.12 | 97.77 | 87.73 | 89.23 | 52.31 | 56.93 |
e-prop | 96.44 | 96.80 | 86.15 | 87.20 | 51.12 | 54.99 |
DECOLLE | 93.61 | 90.52 | 84.16 | 85.17 | 42.30 | 45.70 |
@@
In our investigations, we have observed that the introduction of recurrent weights in the REC architecture yields significant performance improvements when classifying the DVS Gesture and TIMIT datasets, in contrast to the FF architecture. Table III shows SNN performance across various training methods for N-MNIST, DVS Gesture, and TIMIT classification. When trained with BPTT, the REC architecture exhibits a notable enhancement in accuracy when compared to the FF architecture. Similar accuracy improvements are also observed in the case of e-prop and DECOLLE. Benefits of recurrent weights are not evident in the N-MNIST classification task. Here, the performance of SNNs trained with BPTT and e-prop displays only slight differences between the REC module and the FF module. Interestingly, the explicit inclusion of recurrent weights leads to a performance degradation in SNNs trained with DECOLLE for the N-MNIST dataset. Further investigation is required to understand the underlying reason for this degradation.
@@
Additionally, a more local learning method slightly degrades the performance of SNNs, which indicates a trade-off between biological plausibility and performance. In this study, the locality of the learning method is considered to increase from BPTT (global learning method) to e-prop and DECOLLE successively. In N-MNIST classification, a slight accuracy difference is observed between BPTT and e-prop in both REC and FF modules. DECOLLE exhibits an accuracy reduction in both architectures compared with e-prop. A similar trend is observed for DVS Gesture and TIMIT classification tasks. In subsequent sections, we will focus on analyzing the performance differences and robustness between these methods. To provide a case study, we have selected the N-MNIST dataset.
@@ IV-C Representational Differences
CKA is a technique used to measure the similarity between two sets of data, and it has been increasingly applied to understand the internal representation in neural networks. Pioneering work by Cortes et al. introduced CKA as an alignment measure for kernels in the context of support vector machines [52]. Later, Kornblith et al. adapted CKA for deep learning, using it to compare layers across different neural network models, which provided insights into which layers of networks learn similar features [43]. Williams et al. expanded this analysis to study the similarity between representations learned by convolutional neural networks [53]. Li et al. investigated how SNNs differ from traditional ANNs in learning representations, using CKA to analyze both spatial and temporal dimensions [29]. In this study, CKA is employed as a metric to assess the similarity of representations across layers that are learned using various methods with differing degrees of local learning. Consistent with prior works, the CKA is averaged over 4096 instances from the testing set of the N-MNIST dataset [29, 45]. We also compare the similarity of learned representations of the various training methods with ANN-SNN conversion, where the SNN is trained using conventional ANN training methods [54, 55].
@@ IV-D Robustness under Attack
Epsilon | 0.001 | 0.005 | 0.01 | 0.02 | 0.05 |
Training Methods | FF | ||||
BPTT | 83.82 | 16.78 | 4.34 | 1.22 | 0.73 |
e-prop | 89.90 | 41.27 | 12.33 | 3.27 | 2.87 |
DECOLLE | 83.48 | 45.11 | 28.57 | 21.29 | 17.79 |
Training Methods | REC | ||||
BPTT | 94.66 | 80.76 | 58.73 | 27.69 | 18.55 |
e-prop | 93.04 | 78.28 | 60.57 | 35.07 | 18.25 |
DECOLLE | 86.76 | 84.14 | 72.81 | 47.56 | 23.75 |
@@
DECOLLE | 90.21% | 0.86% | 87.46% | 23.75% | 81.55% | 74.10% | 81.29% | 88.96% | 80.81% | 91.02% |
Specifically, the Fast Gradient Sign Method (FGSM) attack leverages the sign of gradients to create adversarial instances that maximize the loss of a neural network, misleading it to make incorrect classifications [56]. However, due to the non-differentiable nature of SNNs, the direct calculation of gradients is not applicable. To solve this problem, previous work has proposed an effective framework which involves crafting an adversarial example for an SNN by creating its ANN counterpart and then calculating gradients for the corresponding ANN [57]. These gradients are subsequently utilized to construct an adversarial example for the SNN. We adopt this variation of the FGSM attack to compare the robustness of various training methods and to investigate the impact of recurrent weights on robustness.
@@
Table IV reports the performance of different training methods when subjected to FGSM attacks, where the perturbation level, denoted as “Epsilon”, varies across a range from 0.001 to 0.05. The results of our benchmarking analysis for backdoor attacks are illustrated in Table V. We mainly focus on poison rates in realistic settings below 50% as shown in prior works [58, 59, 60, 34]. Both attacks are evaluated on the N-MNIST dataset.
Enhancement of the model’s robustness is observed with the inclusion of explicit recurrent weights. Under FGSM attacks, compared to FF models trained by the same methods and subjected to the same “Epsilon” values, the REC model retains more accuracy than the FF model. Besides, the robustness advantage of explicit recurrence is also observed under backdoor attack. Specifically, under more realistic settings with poison rates at 10% and 30%, we observe that the REC model consistently exhibits a lower ASR compared to its corresponding FF model in all three training methods.
@@ IV-E CKA Measurements to Explain Robustness
It is observed that the addition of recurrent weights contributes to better preservation of the representations learned at each layer across these training methods. In these experiments, positive similarity differences are obtained at all perturbation levels, revealing that REC SNNs maintain larger representational similarity (implying more robust representations learned in the presence of attacks) compared to FF SNNs between clean and adversarial data. The results also reveal that global learning methods exhibit a greater enhancement in robustness with recurrent connections, as the similarity differences are larger in BPTT than in e-prop and DECOLLE. The trend matches the data provided in Table IV. Through this analysis, we provide evidence that the inclusion of recurrent weights enhances the robustness of neuromorphic architectures.
@@ V Conclusions
Edge devices like smartphones, Internet of things (IoT) gadgets, and embedded systems require improved energy efficiency due to strict memory and computational constraints [64, 65]. Recent advancements in local learning methods enable on-chip learning [66, 67], allowing models to be trained and updated directly on edge devices with significantly reduced energy consumption compared to global learning methods [68, 15, 19]. In this work, these local learning methods are compared in terms of performance and robustness. To understand the performance disparity of these methods, representational differences are examined through the CKA metric. SNN learning methods with a higher degree of locality are shown to have performance degradation, but they show better robustness under FGSM adversarial attacks. However, the robustness advantages of local learning methods are not observed for non-gradient-based attacks like backdoor attacks. Additionally, we investigate the contribution of additional recurrent weights on the performance of SNNs. Measurements on the relative importance of recurrent weights are performed using Fisher information. We illustrate that recurrent weights not only have positive influences on the performance of SNNs but also augment the robustness of SNNs. These observations are further substantiated by the fact that recurrent weights and linear weights exhibit comparable informative significance. Overall, this work provides valuable insights into the selection of learning methods for SNNs and offers a benchmarking framework for future SNN designers. We acknowledge that our conclusions may currently be limited to feed-forward architectures. Our observations could potentially be generalized to more complex network topologies. Additionally, the integration of more advanced analytical tools could further enhance these insights. Future work could also focus on extending our analysis to convolutional layers [69, 70] and transformer-based SNN models [71, 72, 73].
@@ Acknowledgments
This material is based upon work supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number #DE-SC0021562, the U.S. National Science Foundation under award No. CCF #1955815, CAREER # 2337646 and EFRI BRAID #2318101 and by Oracle Cloud credits and related resources provided by the Oracle for Research program.
Appendix A SNN Accuracy in Convolutional Architectures
In this section, the DVS Gesture dataset is classified by FF and REC SNNs consisting of 3 convolutional layers with 64, 128, and 128 channels respectively (each with a kernel size of 7, a stride of 1, and a padding of 1). The observations (Table VI) are consistent with those in Section IV-B.
Training | DVS Gesture | |
Methods | FF | REC |
BPTT | ||
DECOLLE |
Appendix B Hyperparameter Configuration
Table VII summarizes the hyper-parameters for optimal performance achieved by each method, where LR represents learning rate and BS represents batch size.
Model | LR | BS | Epochs | |||
N-MNIST | ||||||
BP FF | 0.9 | 0.5 | 0.9 | 1e-4 | 16 | 100 |
e-prop FF | 0.99 | 0.95 | 0.2 | 5e-3 | 5 | 100 |
DECOLLE FF | 0.97 | 0.92 | 1.0 | 1e-3 | 72 | 100 |
BP REC | 0.9 | 0.5 | 0.9 | 1e-2 | 256 | 100 |
e-prop REC | 0.99 | 0.95 | 0.2 | 5e-3 | 4 | 100 |
DECOLLE REC | 0.97 | 0.92 | 1.0 | 1e-5 | 72 | 100 |
DVS-Gesture | ||||||
BP FF | 0.9 | 0.5 | 1.0 | 1e-3 | 16 | 100 |
e-prop FF | 0.95 | 0.65 | 0.3 | 1e-3 | 15 | 100 |
DECOLLE FF | 0.9 | 0.65 | 0.9 | 2e-4 | 72 | 100 |
BP REC | 0.95 | 0.5 | 0.9 | 1e-3 | 32 | 100 |
e-prop REC | 0.95 | 0.6 | 0.7 | 3e-3 | 15 | 100 |
DECOLLE REC | 0.05 | 0.2 | 1.0 | 3e-5 | 72 | 100 |
TIMIT | ||||||
BP FF | 0.9 | 0.5 | 1.0 | 1e-3 | 8 | 100 |
e-prop FF | 0.99 | 0.3 | 1.0 | 5e-4 | 4 | 100 |
DECOLLE FF | 0.97 | 0.9 | 0.9 | 2e-4 | 128 | 100 |
BP REC | 0.9 | 0.7 | 1.0 | 1e-3 | 8 | 100 |
e-prop REC | 0.99 | 0.3 | 1.0 | 5e-4 | 4 | 100 |
DECOLLE REC | 0.97 | 0.3 | 1.0 | 1e-5 | 32 | 100 |